Find the amount that results from the investment.
$480 invested at 10% compounded quarterly after a period of 6 years

Answer by mananth(16946)

Principal P = 480
Amount= A
years=n 6.00
compounded 4 times a year t
Rate = 10.00 0.1
Amount = P*((n+r)/n)^n*t

Amount =480*(1+ 0.1/t)^ 6*4
Amount =480*(1+ 0.03)^ 24
480 *( 1.03 )^ 24
Amount = 868.19